# Evaluation and Facilitation of Online Discussions in the LLM Era: A Survey

### v2

@@ Abstract
We present a survey of methods for assessing and enhancing the quality of online discussions, focusing on the potential of Large Language Models (LLMs). While online discourses aim, at least in theory, to foster mutual understanding, they often devolve into harmful exchanges, such as hate speech, threatening social cohesion and democratic values. Recent advancements in LLMs enable artificial facilitation agents to not only moderate content, but also actively improve the quality of interactions. Our survey synthesizes ideas from Natural Language Processing (NLP) and Social Sciences to provide (a) a new taxonomy on discussion quality evaluation, (b) an overview of intervention and facilitation strategies, (c) along with a new taxonomy of conversation facilitation datasets, (d) an Large Language Model (LLM)-oriented roadmap of good practices and future research directions, from technological and societal perspectives.
@@ A Survey
[…]
@@ 1 Introduction
Social Science provides theories and applications for the facilitation of a discussion, but in specific contexts, such as teaching/learning Mansour (2024) or clinical discussions Gelula (1997), without much research devoted to online discussions, such as in social media. While prior NLP studies have explored LLM-facilitated discussions Burton et al. (2024); Aher et al. (2023); Beck et al. (2024); Schroeder et al. (2024); Small et al. (2023); Cho et al. (2024), rarely does Social Science work examine how facilitation can be automated Gimpel et al. (2024).
@@
The main research question of this survey is can LLMs be used effectively as facilitators in online discussions? To explore this question, we focus on three key areas: (1) methods (potentially also LLM-based) for evaluating aspects of online discussions, (2) intervention strategies for facilitation, and (3) available data resources relevant to facilitation. Specifically, we survey discussion evaluation aspects and introduce a new taxonomy (§4). We map tasks suited for ML models, LLMs, and humans, aggregate multidimensional insights on facilitation strategies (§5), and outline future possibilities for LLMs (§6). Additionally, we compare major datasets, dividing them into categories per task (§7). Our work focuses mostly on written thread-like discussions (§2).
@@ 2 Terminology
(or “threads”) are discussions which start from an Original Post (OP) with subsequent comments replying to either the OP or to other comments Seering (2020).
@@ 4 Discussion Quality Evaluation
In this work, we draw from the works of Bächtiger et al. (2022, 2010); Steenbergen et al. (2003); Falk and Lapesa (2023) and Kies (2022) to define a new social-science-informed taxonomy for discussion quality dimensions. While we present a structured taxonomy, it is important to note that the categories are not mutually exclusive. Rather, elements within the taxonomy may coexist within evaluation dimensions, complement one another, or serve as explanatory mechanisms for other dimensions. An example of the dimension interaction can be found in Table F in the Appendix. The grouped dimensions along with the NLP approaches are shown in the Appendix in Table 4.
@@ 4.1 Structure and Logic
‘Coherence’, as described above, evaluates logical consistency, while ‘flow’ assesses smooth progression in discussions Li et al. (2021). Both are essential tools for facilitators in their effort to redirect off-topic comments and guide transitions between topics during a discussion Lambert et al. (2024); Park et al. (2012); Falk et al. (2024). A sudden drop in how well responses match the topic or question often comes before personal attacks or off-topic turns (Chang and Danescu-Niculescu-Mizil, 2019; Zhang et al., 2018), making coherence and flow indicators of argument structure and a valuable early signal for facilitators.
@@
Rooted in Speech Act Theory Austin (1975); Searle (1969), dialogue acts have been employed to assess deliberative quality and analyze facilitation strategies Fournier-Tombs and MacKenzie (2021); Chen et al. (2024a). They characterize dialogue turns (e.g., interruption) to analyze interaction dynamics Ferschke et al. (2012); Stolcke et al. (2000); Zhang et al. (2017); Al-Khatib et al. (2018). Positive (e.g., causal reasoning) or negative (e.g., disrespect) dialogue acts can be scored to reflect discussion quality and low scores may indicate the need for interventions Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024).
@@ 4.2 Social Dynamics
Disagreements, when constructive, improve discussions by fostering deeper understanding Friess (2018); De Kock and Vlachos (2021). Assessing disagreement, however, is complex. The hierarchy of Graham, 2008 considers disagreement tactics ranging from name calling to refuting the central point. Along with other work on dispute tactics Walker et al. (2012); Benesch et al. (2016); De Kock et al. (2022), it can be used to examine types of disagreements in a discussion.
@@ 4.3 Emotion and Behavior
Empathy is the ability to understand others’ perspectives and emotions and respond correspondingly Lipman (2003); Xu and Jiang (2024). Facilitators desire to foster empathy in online discussions, since it encourages prosocial behavior and boosts engagement Xu and Jiang (2024); Concannon and Tomalin (2024); Lambert et al. (2024). To do so, they encourage users to share personal stories and experiences Schroeder et al. (2024). Various coding schemes Macagno et al. (2022), psychological indicators (e.g., the emotion-laden words of Furniturewala and Jaidka, 2024), and dimensions (e.g., perceived engagement such in Xu and Jiang, 2024) have been used to detect both expressed and perceived empathetic traits.
@@
Controversy arises from divergent viewpoints, leading to polarized exchanges that can escalate to toxicity and derail online discussions Avalle et al. (2024). Controversial comments have been shown to contribute to a decline in positive emotions and a sustained rise in anger Hessel and Lee (2019); Chen et al. (2024b). The spread of political leanings among discussants and sentiment distribution analysis are common approaches to measure controversy Avalle et al. (2024).
@@ 4.5 LLM Approaches to Discussion Quality
LLMs can significantly aid in evaluating discussion quality, performing on par with humans in annotating argument structure, coherence, and flow across tasks like argument mining and synthesis Mirzakhmedova et al. (2024); Rescala et al. (2024); Wang et al. (2023); Chen et al. (2024a); Irani et al. (2024); Anastasiou and De Liddo (2024); Zhang et al. (2024); Mendonca et al. (2024); Zhang et al. (2023). LLMs can also reliably label dialogue acts Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024), as well politeness, power, disagreement, and toxicity Zhou et al. (2024); Ziems et al. (2024). However, they struggle with tasks involving social norms (such as irony and humor) and show limited accuracy in sentiment and engagement detection Hu et al. (2023); Sravanthi et al. (2024); Furniturewala and Jaidka (2024); Xu and Jiang (2024). While LLMs show promise in measuring controversy and persuasion, performance drops at the discussion level, especially when aiming to measure diversity, informativeness, and generally when social and pragmatic understanding is necessary Ziems et al. (2024); Avalle et al. (2024); Lawrence and Reed (2020).
@@ 5.1 When to Intervene
Picking the right moment to intervene is a crucial part of effective facilitation strategies. If a facilitator does not intervene when they should have, there is a risk of significant escalation, while intervening when unnecessary can increase toxicity Schaffner et al. (2024); Trujillo and Cresci (2022); Schluger et al. (2022); Cresci et al. (2022). It is imperative then (also considering the evaluation dimensions discussed in §4), for a facilitator to be able to recognize subtle cues hinting towards escalation, in order to defuse the situation, something that even experienced human facilitators are not confident to reliably do Schluger et al. (2022). The NLP task of ‘Conversational Forecasting’ may contribute towards this direction. Given a conversation up to a point, a model attempts to predict if an event will occur in the future in that conversation. In our case, this is where a facilitator would intervene Schluger et al. (2022). Traditional ML models can perform well on this task, although their performance varies Falk et al. (2021); Park et al. (2012); Falk et al. (2024); Schluger et al. (2022).
@@ 5.2 How to Intervene
There is currently no agreed-upon taxonomy for facilitator interventions. Lim et al. (2011) propose a taxonomy that focuses on discussion facilitation, excluding, however, disciplinary or administrative actions, which are common in online discussions. Park et al. (2012) propose another taxonomy consisting of seven moderator functions, ranging from policing the discussion to solving technical issues. These functions roughly correlate with the volunteer moderator roles, as described by Seering (2020). More practical approaches can be found in facilitator manuals eRulemaking Initiative (2017); MIT Center for Constructive Communication (2024) and books White et al. (2024).
Facilitators often have to decide what form of coercive measure to take to make sure the conversation remains healthy, without having to intervene repeatedly. Human interventions typically use an unofficial ‘escalation ladder’ (Figure 1), where the facilitator will progressively move from milder facilitation tactics to threatening, and finally disciplinary action Seering (2020). ‘Conversational moderation’ Cho et al. (2024), where a facilitator first converses with the offender, has proven effective and is actively encouraged in some facilitator guidelines The Commons (2025). This is probably why disciplinary action is typically not the first choice of a facilitator Schluger et al. (2022) and why it should reasonably be used as a last resort. Softer kinds of interventions that facilitators frequently use first include: setting and informing users about rules Schluger et al. (2022); Seering (2020), welcoming new users Schluger et al. (2022), summarizing key points Small et al. (2023); Falk et al. (2024), balancing participation Kim et al. (2021); Fishkin et al. (2018), and aiding users improve their points Tsai et al. (2024); Falk et al. (2024).
@@ 5.3 Personalized Interventions
It is worth stressing that intervention strategies should not be applied en masse, without considering the characteristics of each individual. Traditionally, massive application (or threatening) of disciplinary action has led to adverse effects community- and platform-wide Trujillo and Cresci (2022); Falk et al. (2021) and the creation of echo-chambers Cho et al. (2024). There are also calls for research to move away from one-size-fits-all approaches and instead move towards personalized interventions Cresci et al. (2022). Human facilitators are often able to personalize interventions per individual Schluger et al. (2022), and we hypothesize that LLMs can also do so to some extent.
@@ 6 Towards LLM-based facilitation
Advances in LLMs enable the development of facilitation agents that more actively engage in discussions. These agents can warn users about policy violations Kumar et al. (2024), suggest rephrasings to improve tone or persuasiveness Bose et al. (2023), monitor turn-taking Schroeder et al. (2024), and summarize or visualize key discussion points Small et al. (2023). They can also assist in drafting group statements that reflect diverse viewpoints Tessler et al. (2024). A brief, non-exhaustive summary of the capabilities of simpler ML models, LLMs, and humans can be found in Figure 2.
@@ 6.3 Fully Automated LLM-based Facilitation
There are indications that LLMs can be used as facilitators in the fullest capacity of the role. LLMs are able to predict optimal facilitation tactics Schroeder et al. (2024), like traditional ML models Al-Khatib et al. (2018). Furthermore, they have proven capable of developing and executing social strategies in other tasks, e.g., negotation games, LLM interactions Abdelnabi et al. (2024); Cheng et al. (2024a); Martinenghi et al. (2024). Given that relatively simple ML chatbots, which do not leverage generative text capabilities, have been reported to improve discussions Kim et al. (2021), many expect LLM-based facilitation to be a promising solution to the well-known bottleneck of human facilitation Small et al. (2023); Seering (2020); Burton et al. (2024); Schroeder et al. (2024). Notably, Cho et al. (2024) successfully use LLM facilitators with prompts based on Cognitive Behavioral Therapy to moderate a live discussion with human participants. Their work shows that LLM facilitators can adapt in their instructions to users, although they cannot by themselves affect the discussion with regard to cooperation and mutual respect between the participants.
Nevertheless, LLMs have inherent limitations that make them worse than humans in most social tasks (Figure 2) Rossi et al. (2024). While human facilitators are encouraged to be neutral White et al. (2024); eRulemaking Initiative (2017), numerous studies point to biases in sociodemographic, statistical, and political terms in LLMs R.Anthis et al. (2025); Hewitt et al. (2024); Rossi et al. (2024), which can be exacerbated during the course of a discussion Taubenfeld et al. (2024).

### v3

@@ Abstract
We present a survey of methods for assessing and enhancing the quality of online discussions, focusing on the potential of Large Language Models (LLMs). While online discourses aim, at least in theory, to foster mutual understanding, they often devolve into harmful exchanges, such as hate speech, threatening social cohesion and democratic values. Recent advancements in Large Language Models enable artificial facilitation agents to not only moderate content, but also actively improve the quality of interactions. Our survey synthesizes ideas from Natural Language Processing (NLP) and Social Sciences to provide (a) a new taxonomy on discussion quality evaluation, (b) an overview of intervention and facilitation strategies, (c) along with a new taxonomy of conversation facilitation datasets, (d) an LLM-oriented roadmap of good practices and future research directions, from technological and societal perspectives.
@@ A Survey
Katerina Korre† , Dimitris Tsirmpas†‡, Nikos Gkoumas†, Emma Cabalé♢ Danai Myrtzani‡ , Theodoros Evgeniou♠ Ion Androutsopoulos‡†, John Pavlopoulos‡† †Archimedes, Athena Research Center, Greece ({k.korre,n.goumas}@athenarc.gr) ‡Athens University of Economics and Business, Greece ({dim.tsirmpas,dan.myrtzani,ion,annis}@aueb.gr) ♢École Normale Supérieure Paris-Saclay, France (emma.cabale@ens-paris-saclay.fr) ♠INSEAD, Technology and Business, France (theodoros.evgeniou@insead.edu)
@@ 1 Introduction
Social Science provides theories and applications for the facilitation of a discussion, but in specific contexts, such as teaching Mansour (2024) or clinical discussions Gelula (1997), without much research devoted to online discussions. While prior NLP studies have explored LLM-facilitated discussions Burton et al. (2024); Aher et al. (2023); Beck et al. (2024); Schroeder et al. (2024); Small et al. (2023); Cho et al. (2024), rarely does Social Science work examine how facilitation can be automated Gimpel et al. (2024).
@@
The main research question of this survey is can LLMs be used effectively as facilitators in online discussions? Focusing on threaded discussions (§2), we explore three key areas: (1) methods (potentially also LLM-based) for evaluating aspects of online discussions, (2) intervention strategies for facilitation, and (3) available data resources which can be used to analyze human facilitation and train LLM equivalents. Specifically, we survey discussion evaluation aspects and introduce a new taxonomy (§4). We map tasks suited for ML models, LLMs, and humans, aggregate multidimensional insights on facilitation strategies (§5), and outline future possibilities for LLMs (§6). Additionally, we aggregate and compare all major relevant datasets in literature, dividing them into categories per task (§7).
@@ 2 Terminology
(or ‘threads’) are discussions which start from an Original Post (OP) with subsequent comments replying to either the OP or to other comments Seering (2020).
@@ 4 Discussion Quality Evaluation
In this work, we draw from the works of Bächtiger et al. (2022, 2010); Steenbergen et al. (2003); Falk and Lapesa (2023) and Kies (2022) to define a new social-science-informed taxonomy for discussion quality dimensions. While we present a structured taxonomy, it is important to note that the categories are not mutually exclusive. Rather, elements within the taxonomy may coexist within evaluation dimensions, complement one another, or serve as explanatory mechanisms for other dimensions. An example of the dimension interaction can be found in Table 5 in the Appendix F. The grouped dimensions along with the NLP approaches are shown in the Appendix in Table 4.
@@ 4.1 Structure and Logic
Coherence, as described above, evaluates logical consistency, while flow assesses smooth progression in discussions Li et al. (2021). Both are essential tools for facilitators in their effort to redirect off-topic comments and guide transitions between topics during a discussion Lambert et al. (2024); Park et al. (2012); Falk et al. (2024). A sudden drop in how well responses match the topic or question often comes before personal attacks or off-topic turns (Chang and Danescu-Niculescu-Mizil, 2019; Zhang et al., 2018), making coherence and flow indicators of argument structure and a valuable early signal for facilitators.
@@
Rooted in Speech Act Theory Austin (1975); Searle (1969), dialogue acts have been employed to assess deliberative quality and analyze facilitation strategies Fournier-Tombs and MacKenzie (2021); Chen et al. (2024). They characterize dialogue turns (e.g., interruption) to analyze interaction dynamics Ferschke et al. (2012); Stolcke et al. (2000); Zhang et al. (2017); Al-Khatib et al. (2018). Positive (e.g., causal reasoning) or negative (e.g., disrespect) dialogue acts can be scored to reflect discussion quality with low scores potentially indicating a need for intervention Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024).
@@ 4.2 Social Dynamics
Disagreements, when constructive, improve discussions by fostering deeper understanding Friess (2018); De Kock and Vlachos (2021). Assessing disagreement, however, is complex. The hierarchy of Graham (2008) considers disagreement tactics ranging from name calling to refuting the central point. Along with other work on dispute tactics Walker et al. (2012); Benesch et al. (2016); De Kock et al. (2022), it can be used to examine types of disagreements in a discussion.
@@ 4.3 Emotion and Behavior
Empathy is the ability to understand other perspectives and emotions and respond correspondingly Lipman (2003); Xu and Jiang (2024). Facilitators desire to foster empathy in online discussions, since it encourages prosocial behavior and boosts engagement Xu and Jiang (2024); Concannon and Tomalin (2024); Lambert et al. (2024). To do so, they encourage users to share personal stories and experiences Schroeder et al. (2024). Various coding schemes Macagno et al. (2022), psychological indicators (e.g., the emotion-laden words of Furniturewala and Jaidka, 2024), and dimensions (e.g., perceived engagement such as in Xu and Jiang, 2024) have been used to detect both expressed and perceived empathetic traits.
@@
Controversy arises from divergent viewpoints, leading to polarized exchanges that can escalate to toxicity and derail online discussions Avalle et al. (2024). Controversial comments have been shown to contribute to a decline in positive emotions and a sustained rise in anger Hessel and Lee (2019); Chen et al. (2025). The spread of political leanings among discussants and sentiment distribution analysis are common approaches to measure controversy Avalle et al. (2024).
@@ 4.5 LLM Approaches to Discussion Quality
LLMs can significantly aid in evaluating discussion quality, performing on par with humans in annotating argument structure Mirzakhmedova et al. (2024); Rescala et al. (2024), excelling in comparative argument evaluation Wang et al. (2023), AM, and synthesis Chen et al. (2024); Irani et al. (2024); Anastasiou and De Liddo (2024). They are increasingly used for coherence evaluation at the comment or whole discussion level Zhang et al. (2024), often using proprietary models (e.g., GPT-4), while fine-tuned open-source models also show promise Mendonca et al. (2024); Zhang et al. (2023). LLMs are not preferred for turn-taking or linguistic markers. Research on the former focuses on visual dashboards (such as VisArgue or TurnViz) that reveal dominance shifts at a glance (El-Assady et al., 2017; Hoque and Carenini, 2016), while distinguishing linguistic markers is often approached through older methodologies such as LSTMs Sak et al. (2014) or dictionaries, as mentioned in §4.1.
LLMs can also serve as dialogue and speech act annotators Ziems et al. (2024); Cimino et al. (2024); Martinenghi et al. (2024); Schroeder et al. (2024). For example, Yu et al. (2024) show that GPT-4 reached almost human accuracy on the task of annotating the speech act of apologizing. However, we acknowledge that the difficulty of automatic speech act annotation might depend on the task and more research on that is encouraged.
Remaining on the frontier of pragmatics, research shows that LLM-fine-tuning enhances implicature comprehension Ruis et al. (2023), with GPT-4 achieving human-level performance through chain-of-thought prompting. While LLMs perform well in some pragmatics tasks, such as in the Pragmatic Understanding Benchmark (PUB) Sravanthi et al. (2024), they struggle with social norm-based understanding (e.g., humor, irony) Hu et al. (2023); Sravanthi et al. (2024). This is also true for annotating politeness, power, disagreement, and toxicity Zhou et al. (2024); Ziems et al. (2024).
LLMs perform well in identifying power differentials in discussions Ziems et al. (2024), and can detect these imbalances in real time, enabling facilitators to invite quieter voices and limit dominant turns. Additionally, LLMs have been successfully employed as dispute tactics annotators, highlighting instances of hostile interactions that may require moderator intervention Zhou et al. (2024). However, they show limited accuracy in sentiment and engagement detection Hu et al. (2023); Sravanthi et al. (2024); Furniturewala and Jaidka (2024); Xu and Jiang (2024). Empathy detection also remains challenging for LLMs, with evaluations showing inconsistent performance across conversational tasks Furniturewala and Jaidka (2024); Xu and Jiang (2024); Ziems et al. (2024). While LLMs show promise in measuring controversy and persuasion, performance drops at the discussion level, particularly when assessing diversity, informativeness, and broader aspects of sociopragmatic understanding Ziems et al. (2024); Avalle et al. (2024); Lawrence and Reed (2020).
@@ 5.1 When to Intervene
Picking the right moment to intervene is a crucial part of effective facilitation strategies. If a facilitator does not intervene when they should have, there is a risk of significant escalation, while intervening when unnecessary can increase toxicity Schaffner et al. (2024); Trujillo and Cresci (2022); Schluger et al. (2022); Cresci et al. (2022). Even ‘softer’ interventions such as information and opinion sharing can prove detrimental to discussions when performed excessively Gao et al. (2025). It is imperative then for a facilitator to be able to recognize subtle cues that hint towards escalation (also considering the evaluation dimensions discussed in §4), in order to defuse the situation, something that even experienced human facilitators are not confident to reliably do Schluger et al. (2022).
The NLP task of ‘Conversational Forecasting’ may contribute towards this direction. Given a conversation up to a point, a model attempts to predict if an event will occur in the future in that conversation. In our case, this is where a facilitator would intervene Schluger et al. (2022). Traditional ML models can perform well on this task, although their performance varies Falk et al. (2021); Park et al. (2012); Falk et al. (2024); Schluger et al. (2022).
@@ 5.2 How to Intervene
There is currently no agreed-upon taxonomy for facilitator interventions. Lim et al. (2011) propose a taxonomy that focuses on discussion facilitation, excluding, however, disciplinary or administrative actions, which are common in online discussions. Park et al. (2012) propose another taxonomy consisting of seven moderator functions, ranging from policing the discussion to solving technical issues. Their taxonomy, however, is not easily generalizable to domains other than website facilitation Chen et al. (2025). These functions roughly correlate with the volunteer moderator roles, as described by Seering (2020). More practical approaches can be found in facilitator manuals eRulemaking Initiative (2017); MIT Center for Constructive Communication (2024) and books White et al. (2024). Chen et al. (2025) bridge the questions of when and how to facilitate by proposing a taxonomy that analyzes both individually, which was improved by Gao et al. (2025).
Several works have examined facilitation in education. Sjølie et al. (2021) conducted a mixed-methods study on a meta-facilitative approach, where students and teachers explicitly discussed their collaboration and which led to significant learning improvements. In the context of virtual facilitation, Verkuyl et al. (2024) showed that successful integration of virtual simulations in higher education depends not just on access, but on facilitators who align simulations with course objectives, respond to learners’ needs, and evaluate the experience. Both studies suggest that facilitation requires socially informed practices, even as automation promises workload reduction.
With reference to NLP approaches in facilitation in education, Lugini et al. (2020) designed Discussion Tracker, a classroom analytics tool that applies algorithms to identify argumentation moves (claim, evidence, explanation) and evaluate levels of specificity, as well as recognize patterns of collaboration. Deployment in class showed that teachers considered the analytics valuable, and that the system’s classifiers achieved moderate to substantial agreement with human judgments. The aforementioned work of Gao et al. (2025) presented an approach that combines automatic English as a Second Language (ESL) dialogue assessment with a framework of moderation strategies. The authors showed that moderators improve topic flow and conversation management, with active acknowledgment and encouragement proving most effective, but excessive input can hinder discussion.
Facilitators often have to decide what form of coercive measure to take to make sure the conversation remains healthy, without having to intervene repeatedly. Human interventions typically use an unofficial ‘escalation ladder’ (Figure 1), where the facilitator will progressively move from milder facilitation tactics to threatening, and finally enacting disciplinary action Seering (2020). ‘Conversational moderation’ Cho et al. (2024), where a facilitator first converses with the offender, has proven effective and is actively encouraged in some facilitator guidelines The Commons (2025). This is probably why disciplinary action is typically not the first choice of a facilitator Schluger et al. (2022) and why it should reasonably be used as a last resort.
Softer kinds of interventions that facilitators frequently use first include: setting and informing users about rules Schluger et al. (2022); Seering (2020), welcoming new users Schluger et al. (2022), summarizing key points Small et al. (2023); Falk et al. (2024), balancing participation Kim et al. (2021); Fishkin et al. (2018), and aiding users improve their points Tsai et al. (2024); Falk et al. (2024). Facilitators are also instrumental in beginning and ending discussions Small et al. (2023); Gao et al. (2025), as well as generally encouraging participants Gao et al. (2025). It is worth noting that facilitation guides may explicitly forbid facilitators from intervening in certain ways, such as sharing their opinions or providing new information MIT Center for Constructive Communication (2024).
@@ 5.3 Personalized Interventions
Intervention strategies should not be applied en masse, without considering the characteristics of each individual. Traditionally, massive application of disciplinary action (or threatening) has led to adverse effects community- and platform-wide Trujillo and Cresci (2022); Falk et al. (2021) and to the creation of echo-chambers Cho et al. (2024). There are also calls for research to move away from one-size-fits-all approaches and instead move towards personalized interventions Cresci et al. (2022). Human facilitators are often able to personalize interventions per individual Schluger et al. (2022), and we hypothesize that LLMs can also do so to some extent.
@@ 6 Towards LLM-based facilitation
Advances in LLMs enable the development of facilitation agents that engage more actively in discussions. These agents can warn users about policy violations Kumar et al. (2024), suggest rephrasings to improve tone or persuasiveness Bose et al. (2023), monitor turn-taking Schroeder et al. (2024), and summarize or visualize key discussion points Small et al. (2023). They can also assist in drafting group statements that reflect diverse viewpoints Tessler et al. (2024). A brief, non-exhaustive summary of the capabilities of simpler ML models, LLMs, and humans can be found in Figure 2.
@@ 6.3 Fully Automated LLM-based Facilitation
There are indications that LLMs can be used as facilitators to the fullest capacity of the role. LLMs are able to predict optimal facilitation tactics Schroeder et al. (2024), like traditional ML models Al-Khatib et al. (2018). Furthermore, they have proven capable of developing and executing social strategies in other tasks, e.g., negotation games, LLM interactions Abdelnabi et al. (2024); Cheng et al. (2024a); Martinenghi et al. (2024). Given that relatively simple ML chatbots, which do not leverage generative text capabilities, have been reported to improve discussions Kim et al. (2021), many expect LLM-based facilitation to be a promising solution to the well-known bottleneck of human facilitation Small et al. (2023); Seering (2020); Burton et al. (2024); Schroeder et al. (2024). Notably, Cho et al. (2024) successfully use LLM facilitators with prompts based on Cognitive Behavioral Therapy to moderate a live discussion with human participants. Their work shows that LLM facilitators can adapt their instructions to users, although they cannot by themselves affect the discussion with regard to cooperation and mutual respect between the participants.
Nevertheless, LLMs have inherent limitations that make them worse than humans in most social tasks (Figure 2; Rossi et al. (2024)). While human facilitators are encouraged to be neutral White et al. (2024); eRulemaking Initiative (2017), numerous studies point to biases in sociodemographic, statistical, and political terms in LLMs Anthis et al. (2025); Hewitt et al. (2024); Rossi et al. (2024), which can be exacerbated during the course of a discussion Taubenfeld et al. (2024).